2 3. Find all eigenvalues and eigenvectors of the matrix -2 0 -1 0 2 0 1 -2
Answer #1

Let 2 A = -2 -1 2 O o 1 0 -2 Consider the characteristics equation of A det (\I-A) 0 [..+2 O det det 1 -2 -1 X+2 =O - 2 0 Exc☆ -21, +262 = 0 = -X, +13 - 0 7 x = x3 Also, and - 22, +33 = 0 I, = 23 2354-2219=0 = Oxi: 23 So, Eigenspace, i.eu, space of e2 2 xi -1 4 1 =0 2 22 X3 o 2 → 23, + 2x3 = 0 = 2,= -013 (**) -1, +432 + 73 = 0 7 X₂ + 4x + 2 = 0 (from **) » 2x2 +4866 = * x3